# Timing Matters: Enhancing User Experience through Temporal Prediction in Smart Homes

### v1

@@ Introduction
96-class 8-class 8-class 8-class 8-class 8-class
MLP 0.137 0.682 0.29 0.293 0.395 0.373
LSTM 0.222 0.768 0.34 0.368 0.427 0.442
TCN 0.014 0.199 0.18 0.193 0.212 0.135
Trans + TCN 0.29 0.77 0.34 0.339 0.404 0.432
Timing-Matters 0.383 0.800 0.375 0.368 0.404 0.444
@@
96-class 8-class 96-class
Timing-Matters 0.383 0.800 0.375
Timing-Matters - RBF 0.381 0.789 0.35
Timing-Matters - Time Encoder 0.335 0.782 0.368
Timing-Matters - Sequence Encoder 0.283 0.781 0.372
@@
This dataset is more detailed than the Smartsense Datasets as the AN dataset contains the entire time and date information for each action. The data is represented as a tensor of size $[N\times 10\times 5]$, where N denotes the number of instances. Each instance consists of a sequence of $10$ consecutive actions performed, and each action consists of a vector of length 5, indicating the day of the year, time of day, device, user id and device control. In this context, the time is an integer denoting the number of seconds the action was performed after the start of the day. Here the date refers to an integer denoting the day of the year a particular action was conducted on.
The user id is a novel component in our dataset whereby instead of randomly generating sequences from a multitude of users as was done in the SmartSense datasets, we make sure each generated sequence is marked by which user executes the sequence. We have 39 different users each of whom follow a different pattern when it comes to utilization of smart devices.
Baselines
@@
  • MLP MultiLayer Perceptron is a simple neural network that just flattens all input and utilizes the following 3 hidden layers of size 1024, 128, and finally num_classes.
  • LSTM Long short-term memory consists of stacking together the vectors of each action and using LSTM with hidden layer size $40$ and 2 layers to make predictions.
  • TCN Refers to a simple Temporal Convolution Network structure (See Section 4.2) where we utilize a two-layer TCN with a kernel size of 2 in both the encoder and decoder layers followed by a Linear layer.
  • Trans + TCN: Utilizing two encoders, where Trans is a Transformer Encoder with two heads stacked using two layers followed by a linear layer to encode information about an action into a single vector. This sequence of action vectors is then passed through a Temporal Convolution Network of 2 layers in both the encoder and decoder with a kernel size of 2 followed by a final Linear layer.
Evaluation Metrics
In our problem, time prediction is a classification problem where we have to make our predictions regarding what time interval the next action is likely to be taken. Hence, our final output is the list of probability values for all classes. We rely on two evaluation metrics, 96-class Accuracy and 8-class Accuracy which, if we divide a day into some $k$ number of classes, denote the fraction where our prediction and true values belong to the same class. In our cases, we rely on $k=96$ and $8$ which correspond to classes of time intervals of 15 min and 3 hrs respectively.
Figure 4: Heatmap of frequencies of top devices in AN and FR dataset.
Experimental Process
@@
HyperParameters
The dimension of the vector embedding is 50. All Transformer Encoders used contain multi-head structures with 2 heads and 2 layers of encoders stacked together. The feedforward network hidden layer size is 200. The Temporal Convolution Networks consist of an encoder and a decoder, where both of them contain two stacked 1D convolution units with a kernel size of 2 and a stride of 1. The first linear layer at the end of the Sequence Encoder contains a hidden layer of size 100. For training we rely on a batch size of $64$ with a learning rate of $0.0001$ and l2reg $0.0001$.
Results
Data Analysis
@@ Results
Frequency of Actions in sequence
We used the SmartSense Dataset where each element is a sequence of 10 actions. Similarly in constructing the AN dataset we also create sequences of 10 actions. In such cases, we need to understand the distribution of the actions in a given sequence and more importantly what devices they refer to. Taking a look at the distribution of the counts of the two most frequent devices in the sequence (4) we observe that our synthetically generated dataset’s actions in a sequence very closely match with that of the SmartSense Dataset.
Dataset Models Precision(96) Precision(08) RMSE
FR Regression 0.1445 0.1445 63978.875
Classification 0.3567 0.3567 30225.904
AN Regression 0.2684 0.9329 8894.975
Classification 0.3865 0.7943 8804.339
Table 4: Comparison of Regression and Classification Models
@@
Table 5: Precision values for 2-layer and 4-layer models across different context windows
Time Difference between Consecutive Actions
@@
For the smart home dataset, unlike traditional time series datasets, each action does not occur at specific time intervals. In such cases the consecutive time differences actually can be taken to be a time series where the $k^{th}$ element denotes the time between $k^{th}$ action and $(k+1)^{th}$. We analyze this distribution of time difference between consecutive actions. On comparing (5) we find that SmartSense dataset is more discrete while our AN dataset provides a more continous spectrum.
Regression vs Classification
Time is a continuous variable and hence time prediction is considered to be a regression problem. However in our case it has been seen that regression is not the optimal solution. Instead we divide the day into 96 intervals and predict which interval the time of next action would lie in. Also for the regression, the metric Precision(96) means whether the prediction is within a 15 minute range from true value. As seen in Table 4, both for the Precision(96) and Test RMSE the classification method yields much better results compared to regression. Even though for the AN dataset, Precision(08) is much better, Precision(96) which gives accuracy upto 15 minutes is more better.
Accuracy of Model
@@
We measure the accuracy of the model and the baselines using four SmartSense Datasets and 1 Anonymous dataset. We show both the results in terms of 96 classes and 8 class classifications. Note that the SmartSense datasets have both 96-class and 8-class results, as the datasets themselves contain time information only up to a precision of 3-hour intervals. Table 2 shows that our model outperforms nearly all datasets.
Ablation Study
We take a look at the components of the model and check which ones play an important role in giving us our results (Table 3). We first replace the Radial Basis Functional Embedding with the Time2Vec Embedding to get Timing-Matters - RBF. This ensures we only take into account periodic variations and not radial distance information. Similarly, by replacing the Time Encoder with an Identity function, we end up with Timing-Matters - Time Encoder. Instead of having information about the neighbourhood at each Time Difference element, we have just the element itself.
As the most crucial link, by using Timing-Matters - Sequence Encoder, we replace the Transformer Encoder with an Identity function, thereby ensuring that no correlation is achieved between each action. We compare all 3 of these variants with the original and find out that Timing-Matters gives the highest results, showing that all of the parts are necessary for our model.
Context Window
A larger number of historical sequences allows the model to consider more historical information. However, too much historical information may introduce more uncertainty and not contribute to performance improvement. Table 5 shows that 20 historical sequences enables us to achieve the optimal performance. However due to limitations of the SmartSense dataset, we have demonstrated our model performance for 10 historical sequences.
Conclusions and Future Work
In this paper, we approach a completely new problem related to User Behaviour in Smart Homes where we learn to predict the time at which the user will perform the next action. This is of extreme importance because whenever you want to model a user behaviour, the temporal context plays a very important role. Our model gives the state-of-the-art results on multiple datasets. It learns the temporal patterns very well thereby increasing its accuracy upto 9.3% as compared to its competitors. We also show that how each component of the model is an essential part of its working through Ablation Study.
@@
Developing real-time systems that continuously learn and adapt to user behavior can ensure that predictions remain accurate and relevant. By focusing on both the type and timing of user actions, smart home systems can become more intuitive and responsive, significantly enhancing the user experience.

### v2

@@ 5.1.1 SmartSense Dataset
Precision(96) Precision(08) Precision(08) Precision(08) Precision(08) Precision(08)
MLP 0.0599 0.4314 0.2891 0.2331 0.3171 0.2633
MLP-2Step 0.0365 0.4115 0.2448 0.2201 0.2450 0.2171
LSTM 0.1276 0.6988 0.2656 0.2220 0.2442 0.2216
MLP-LSTM 0.0156 0.1727 0.2344 0.2109 0.2439 0.2196
LSTM-2Step 0.0156 0.1727 0.3151 0.2493 0.2716 0.2820
Transformer [26] 0.0182 0.3316 0.2318 0.1986 0.0867 0.3172
BERT4Rec [23] 0.0269 0.2231 0.2839 0.2305 0.2563 0.3128
CA-RNN [17] 0.1128 0.6658 0.3646 0.3099 0.4014 0.4235
SASRec [9] 0.2318 0.7639 0.3073 0.2311 0.2729 0.2412
SmartSense [8] 0.3247 0.7726 0.3281 0.2832 0.4093 0.3042
Timing-Matters 0.3830 0.8000 0.3750 0.3680 0.4040 0.4440
@@
Precision(96) Precision(08) Precision(08)
TimingMatters 0.383 0.800 0.375
TimingMatters - RBF 0.381 0.789 0.35
TimingMatters - Time Encoder 0.335 0.782 0.368
TimingMatters - Sequence Encoder 0.283 0.781 0.372
@@ 5.1.2 Anonymous (AN) Dataset
This dataset is more detailed than the Smartsense Datasets as the AN dataset contains the entire time and date information for each action. The data is represented as a tensor of size $N\times 10\times 5]$, where N denotes the number of instances. Each instance consists of a sequence of $10$ consecutive actions performed, and each action consists of a vector of length 5, indicating the day of the year, time of day, device, user ID and device control. In this context, the time is an integer denoting the number of seconds the action was performed after the start of the day. Here the date refers to an integer denoting the day of the year a particular action was conducted on.
The user ID is a novel component in our dataset whereby instead of randomly generating sequences from a multitude of users as was done in the SmartSense datasets, we make sure each generated sequence is marked by which user executes the sequence. We have 39 different users each of whom follows a different pattern when it comes to utilization of smart devices.
5.2 Baselines
@@
  • MLP Concatenates all action embeddings after flattening and passes them through 4 fully connected layers to make the final prediction.
  • MLP-2Step Each action is first processed independently through an MLP. The resulting representations are then concatenated and passed through another MLP for prediction.
  • LSTM Concatenates action embeddings at each timestep and feeds the sequence into a 2-layer LSTM to capture temporal dependencies.
  • MLP-LSTM Each action is passed through an MLP before being fed into a 2-layer LSTM for temporal modeling and final prediction.
  • LSTM-2Step Applies an LSTM individually to each action type. The resulting sequences are then passed through another LSTM for final prediction.
  • Transformer [26] Concatenates action embeddings at each timestep and passes them through stacked Transformer encoder layers to capture global sequence patterns.
  • SASRec [9] leverages self-attention to capture long- and short-term dependencies in user action histories, efficiently identifying relevant past actions for predictions
  • BERT4Rec [23] Uses a bidirectional Transformer BERT architecture for recommendation
  • CA-RNN [17] considers context-dependent features uses context-specific transition matrix in RNN cell in a sequential recommendation.
  • SmartSense [8] applies two levels of attention. The model also transfers commonsense knowledge from routine data to better capture user intentions.
5.3 Evaluation Metrics
In our problem, time prediction is a classification problem where we have to make our predictions regarding what time interval the next action is likely to be taken. Hence, our final output is the list of probability values for all classes. We rely on two evaluation metrics, 96-class Accuracy (called Precision(96)) and 8-class (called Precision(08)) which, if we divide a day into some $k$ number of classes, denote the fraction where our prediction and true values belong to the same class. In our cases, we rely on $k=96$ and $8$ which correspond to classes of time intervals of 15 min and 3 hrs respectively.
Figure 4: Heatmap of frequencies of top devices in AN and FR dataset
5.4 Experimental Process
@@
5.5 HyperParameters
The dimension of the vector embedding is 50. All Transformer Encoders used contain multi-head structures with 2 heads and 2 layers of encoders stacked together. The feedforward network hidden layer size is 200. The Temporal Convolution Networks consist of an encoder and a decoder, where both of them contain two stacked 1D convolution units with a kernel size of 2 and a stride of 1. The first linear layer at the end of the Sequence Encoder contains a hidden layer of size 100. For training, we rely on a batch size of $64$ with a learning rate of $0.0001$ and l2reg $0.0001$.
6 Results and Analysis
6.1 Data Analysis
@@
6.1.1 Frequency of Actions in Sequence
We used the SmartSense Dataset where each element is a sequence of 10 actions. Similarly, in constructing the AN dataset we also create sequences of 10 actions. In such cases, we need to understand the distribution of the actions in a given sequence and, more importantly, what devices they refer to. Taking a look at the distribution of the counts of the two most frequent devices in the sequence (Fig. 4), we observe that our synthetically generated dataset’s actions in a sequence very closely match with that of the SmartSense Dataset.
@@ 6.1.1 Frequency of Actions in Sequence
Table 5: Precision values for models with 2 and 4 transformer layers across different context windows for the AN Dataset
6.1.2 Time Difference between Consecutive Actions
@@
For the smart home dataset, unlike traditional time series datasets, each action does not occur at specific time intervals. In such cases, the consecutive time differences can actually be taken to be a time series where the $k^{th}$ element denotes the time between $k^{th}$ action and $(k+1)^{th}$. We analyze this distribution of time difference between consecutive actions. On comparing (Fig. 5) we find that the SmartSense dataset is more discrete, while our AN dataset provides a more continuous spectrum.
6.2 Regression vs Classification
Time is inherently a continuous variable, which naturally suggests treating time prediction as a regression problem. However, in our study, we found that regression is not the most effective approach for our use case. Instead, we opted to divide the day into 96 intervals, predicting which interval the next action’s time would fall into.
To evaluate the regression performance, we use the Precision(96) metric, which checks whether the predicted time falls within a 15-minute range of the true value. As shown in Table 2, the classification approach significantly outperforms regression in both Precision(96) and Test RMSE metrics. Even though the AN dataset shows better results with the coarser Precision(08) (three-hour window), the finer Precision(96) metric, offering accuracy within 15 minutes, is more effective and demonstrates the superiority of the classification method in capturing time predictions with higher precision.
6.3 Accuracy of Model
@@
We measure the accuracy of the model and the baselines using four SmartSense Datasets and 1 Anonymous dataset. We show the AN results in terms of 96-class and 8-class classification precision. Note that the SmartSense datasets have only 8-class results, as the datasets themselves contain time information only up to a precision of 3-hour intervals.
As shown in Table 3, our proposed Timing-Matters model consistently outperforms the baselines across multiple datasets. For the AN dataset, our model achieves a Precision(96) of 0.383, which is a 6% improvement compared to the best-performing baseline. In the SmartSense datasets except KR (FR, SP, and US), our model achieves the highest Precision(08) scores, with improvements ranging from 1% to 6% over the baselines.
Overall, the results demonstrate that our model offers significant performance gains in time prediction tasks, particularly in scenarios requiring finer time granularity.
6.4 Ablation Study
To assess the importance of each component in our Timing-Matters model, we perform an ablation study by selectively replacing key modules and observing the impact on performance (Table 4).
6.4.1 TimingMatters-RBF
We replace the Radial Basis Function (RBF) embedding with the Time2Vec embedding. This adjustment captures only periodic variations while ignoring radial distance information. The results show a slight drop in performance, with Precision(96) for the AN dataset decreasing from 0.383 to 0.381, and Precision(08) dropping from 0.800 to 0.789. Precision(08) decreases for FR dataset too.
6.4.2 TimingMatters-Time Encoder
We replace the Time Encoder with an Identity function, resulting in the loss of temporal neighborhood information for each time difference element. Instead of encoding surrounding context, only the raw time difference is used. This change further reduces performance, with Precision(96) for the AN dataset decreasing to 0.335 and Precision(08) to 0.782. Precision(08) decreases slightly for FR dataset too.
6.4.3 TimingMatters-Sequence Encoder
We replace the Transformer Encoder with an Identity function, effectively removing any correlation between consecutive actions. This modification leads to the most significant performance drop, with Precision(96) for the AN dataset decreasing to 0.283 and Precision(08) to 0.781. Precision(08) decreases slightly for FR dataset too.
6.5 Context Window
Increasing the number of historical sequences allows the model to leverage more past information, but excessive context can lead to noise and diminishing returns. As shown in Table 5, we ran Timing-Matters on the AN dataset and found using 10 historical sequences yields the best results, closely followed by 5. Due to limitations in the SmartSense dataset, we could not study the optimal context window.
6.6 Effect of Bin Size
The number of bins determines the temporal resolution of user activity modeling. Fewer bins simplify prediction but may miss fine-grained patterns, while more bins provide finer granularity at the cost of sparsity and performance. We studied this phenomenon by running Timing-Matters on the AN dataset as it is the only dataset with precise timestamps. As shown in Table 6, a bin size of 96 (15-min interval) strikes an optimal balance, achieving the lowest RMSE. Precision(num_bins) is not used to make the choice due to its definition changing with the number of bins.
Number of Bins 8 12 24 48 96 288
Precision (num_bins) 0.75 0.62 0.48 0.42 0.38 0.14
RMSE 12658 13407 12172 10896 10865 11320
Table 6: Effect of Bin Size on Test Precision and RMSE for the AN Dataset
7 Conclusions and Future Work
In this paper, we addressed the novel challenge of predicting the timing of users’ next actions in smart home environments, a critical aspect for enhancing proactive AI systems. Recognizing the scarcity of suitable public datasets, we introduced a synthesized dataset with fine-grained timestamped action sequences. Our proposed model Timing-Matters achieves the state-of-the-art performance on multiple datasets. It learns the temporal patterns very well, thus increasing its accuracy by 6% compared to the baselines. Ablation studies confirmed the integral contribution of each model component. Furthermore, our comprehensive analyses revealed that: (1) a context window of 10 historical actions optimally balances information gain against noise; (2) discretizing the day into 96 15-minute intervals (bins) provides the best trade-off between temporal resolution and predictive performance; and (3) framing the problem as a classification task significantly outperforms regression-based approaches in achieving precise timing predictions.
@@ 7 Conclusions and Future Work
{ack}
By using the ack environment to insert your (optional) acknowledgements, you can ensure that the text is suppressed whenever you use the doubleblind option. In the final version, acknowledgements may be included on the extra page intended for references.